An Ad for Cold-Ex claims an 80% end-of treatment clinical success rate for the treatment of the flu. Twelve patients are given Cold-Ex and then later checked to see if the treatment was successful. Let's assume the claim is correct:

a) What is the expected number of patients who are successfully treated?

b) What is the standard deviation?
Answer #1

a) What is the expected number of patients who are successfully treated?

We are given

n = 12

p = 0.80

Expected number of patients = n*p = 12*0.8 = 9.6

Expected number of patients = 9.6

b) What is the standard deviation?

Standard deviation = sqrt(npq)

Where, q = 1 – p = 1 – 0.80 = 0.20

We have n = 12, p = 0.80, q = 0.20

Standard deviation = sqrt(12*0.80*0.20)

Standard deviation = 1.385641